In the US, 60 percent of eligible voters vote in presidential election years. For a random sample of 8 eligible voters in the US, find the probability that all of them voted in the last presidential election.
Answer #1

Answer

we have probability p = 60% = 60/100 = 0.60

sample size is n = 8

Using the binomial probability formula

Probability that all of them voted = P(x=8)

we have the formula P(x=r) = inom{n}{r}*p^r*(1-p)^{n-r}

where n = 8, r = 8 and p = 0.60

setting the given values, we get

P(z = 8) = 0.68 * (1-0.6)8-8 = (81/(8-8)! * 8!)) * 0.68 * 0.40 8!/(8 - 8)! 8! 0.6 0.4

this gives us

P(x=8) = (8!/((8-8)!*8!))*0.6^8*0.4^0 = 0.0168

So, required probability is 0.0168